OpenStudy (anonymous):

Two numbers are in the ratio 5:6.If the sum of the numbers is 66,find the larger number

OpenStudy (anonymous):

a/b=5/6 k(a+b) = 66 ______________ k(11)=66 => k=6 so: a=30 b=36